a student needs to prepare 50.0 ml of 1.00m aqueous h2o2 solution calculate the volume of 4.6 mh2o2 stock solution that should be used to prepare the solution


1
Expert's answer
2020-12-01T09:09:26-0500

Proportion 1:

1 mol – 1000 mL

n mol – 50.0 mL

"n = \frac{1 \times 50}{1000} = 0.05 \; mol"

C(H2O2) = 4.6 M

Proportion 2:

4.6 mol – 1000 mL

0.05 mol – y

"y = \frac{0.05 \times 1000}{4.6} = 10.87 \;mL"

Answer: 10.87 mL of 4.6 M H2O2 stock solution should be used to prepare 50.0 mL of 1.0 M solution.